OpenStudy (anonymous):

find the radius of a circle with an area of 66.5cm^2

OpenStudy (anonymous):

A = pi * r^2 66.5 = pi * r^2 66.5 / pi = r^2 r is around 4.60082682 cm.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Become a fan!